### v2

# Frank–Wolfe algorithm for DC optimization problem

Abstract
We propose two versions of the Frank–Wolfe (conditional gradient) algorithm with adaptive step sizes for DC optimization problems, where the objective is the difference of a differentiable convex function with Lipschitz continuous gradient and a convex (possibly non-differentiable) function. The second version builds on the first, using finite differences to approximate the gradient of the first component. Unlike previous approaches relying on curvature or Lipschitz constants, our stepsizes require no such parameters. We also introduce the class of piecewise-star-convex functions, which may be nonconvex and non–differentiable, and show that the first version achieves an $1/k$ convergence rate for minimizing such functions, while ensuring stationarity of all limit points. In the second version, the finite–difference scheme yields relative rather than absolute gradient errors, simplifying the analysis; we prove that all limit points are stationary and that the duality gap decreases at rate $1/\sqrt{k}$.
Keywords: Frank-Wolfe method; DC optimization problem; finite difference; weak-star-convex function.
AMS subject classification: 90C25, 90C60, 90C30, 65K05.
1 Introduction
The DC optimization method involves minimizing a function that can be represented as the difference between two convex functions. We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable. To the best of our knowledge, DC optimization can be traced back to pioneering works such as [42, 43], which introduced the first algorithms for this problem. Since then, the DC optimization has attracted the attention of the mathematical programming community (see for example [7, 11, 13, 32, 44]), not only for its own sake but also because it is an abstract model for several families of practical optimization problems. Although we are not concerned with practical issues at this time, we emphasize that practical applications emerge whenever the natural structure of the problem is modelled as a DC optimization problem, such as sparse generalized eigenvalue problems [41], sparse optimization problems [19], facility location and clustering problems [36].
The Frank-Wolfe algorithm has a long history, dating back to Frank and Wolfe’s work in the 1950s to minimize convex quadratic functions over compact polyhedral sets, see [15]. This method was generalized about ten years later to minimize convex differentiable functions with Lipschitz continuous gradients and compact constraint convex sets, see [33]. Since then, this method has attracted the attention of several researchers who work with continuous optimization, thus becoming also known as the conditional gradient method. One of the factors that explain the interest in this method is its simplicity and ease of implementation. In fact, each method iteration only requires access to a linear minimization oracle over a compact convex set. It is also worth noting that, due to the method’s simplicity, it allows for low storage costs and ready exploration of separability and sparsity, making its application in large-scale problems quite attractive. It is interesting to note that the popularity of this method has increased significantly in recent years as a result of the emergence of several applications in machine learning, see [25, 29, 30]. For all of these reasons, several variants of this method have arisen throughout the years and new properties of it have been discovered, resulting in a large literature on it, papers dealing with this method include [4, 8, 9, 17, 18, 21, 27, 31, 34].
In the present paper, we formulate two versions of Frank–Wolfe algorithm or conditional gradient method to solve DC optimization problem with an adaptive step size. The second version is based on the first and employs finite differences to approximate the gradient $\nabla g$. It is worth mentioning that the Frank-Wolfe algorithm has previously been formulated in the context of DC programming, see [26]. See also [46] which establishes theoretical connections between DC algorithms and convex-concave procedures with the Frank-Wolfe algorithm. In contrast to the previous study, which assume that the curvature/Lipschitz-type constant of f=ghf=g-h is bounded from above, the analysis of the Frank-Wolfe method done here just assumes that the gradient of the first component $g$ of $f$ is Lipschitz continuous. While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous in both formulations, we will not compute the step size using the Lipschitz constant of $\nabla g$, as in earlier formulations that employed the curvature/Lipschitz-type constant of the objective function $f$. The step size will be computed adaptively, based on an idea introduced in [3] (see also [5, 39]) that approximates the Lipschitz constant. It has been shown in previous works that Frank-Wolfe algorithm produces a stationary point with a convergence rate of only ${\cal O}(1/\sqrt{k})$ for non-convex objective functions, see [28] (see also [26]). We introduce the concept of weak-star-convex function notion, which generalizes the star-convex function concept proposed in [37] as well as a few other related concepts, such as [22, 45]. The weak-star-convex objective functions are also taken into account in the convergence analysis of the first method presented here. The convergence rate is $\mathcal{O}(1/k)$ for both the function values and the duality gap, even though piecewise-star-convexs are generally non-differentiable and non-convex. Consequently, the class of functions for which the Frank–Wolfe method achieves an $\mathcal{O}(1/k)$ convergence rate includes piecewise-star-convexs that are differences of convex functions, with the first component having a Lipschitz continuous gradient. This class complements the set of star–convex functions, which was recently shown in [35] to enjoy the same convergence rate. As mentioned previously, the second version of the Frank-Wolfe algorithm uses finite difference to approximate the gradient of the function $g$. Similar to the strategies adopted in [20], the finite difference utilized here to approximate the gradient is computed with the step-size updated adaptively utilizing two previous iterations. Finite difference is an old concept that has been used to approximate derivatives in optimization settings (for example, see [38, Section 8.1]) and has already appeared in the study of the Frank-Wolfe algorithm (for example, see [14, 16, 25, 40]). It is worth noting that all previous applications of finite difference in the Frank-Wolfe algorithm provided approximate gradients with absolute error, whereas the one used here has the advantage of providing us with a relative error and thus being quite simple to analyze. We show that all limit points of the generated sequence for the second version of the Frank-Wolfe algorithm are stationary points for the problem under consideration, and we establish that the rate of convergence is ${\cal O}(1/\sqrt{k})$ for the duality gap.
The remainder of the paper is organized as follows. Section 2 introduces the notation and collects auxiliary results. In Section 3 we define the concept of piecewise–star–convex, present its main properties, and provide illustrative examples. Section 4 states the optimization problem, the required assumptions, and related notation. Section 5 is devoted to the first version of the Frank–Wolfe algorithm, including its formulation, well–posedness, and analysis of the generated sequence. Section 6 presents the second version of the algorithm together with its formulation, well–posedness, and convergence analysis. Finally, Section 7 summarizes our conclusions.
2 Preliminaries
In this section, we recall some notations, definitions and basics results used throughout the paper. A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be convex if $\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)$, for all $x,y\in{\mathbb{R}^{n}}$ and $\lambda\in[0,1]$. And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex function see [23]. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz if, for all $x\in\mathbb{R}^{n}$, there exist a constant ${\cal C}_{x}>0$ and a neighborhood $U_{x}$ of $x$ such that $|f(x)-f(y)|\leq{\cal C}_{x}\|x-y\|$, for all $y\in U_{x}.$ It is well known that, if $f:\mathbb{R}^{n}\to\mathbb{R}$ is convex, then $f$ is locally Lipschitz. And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [10, p. 32]. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial_{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
$f^{\circ}(x;d)=\limsup_{\tiny{\begin{array}[]{c}u\rightarrow x\\ t\downarrow 0\end{array}}}\frac{f(u+td)-f(u)}{t}.$
For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [10, p. 27]. If $f$ is convex, then $\partial_{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [10, p. 36]. We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial_{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [10, p. 33].
Theorem 2.1 ([10, p. 27]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. Then, $\partial_{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial_{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$. Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial_{c}f(x)\}$.
Theorem 2.2 ([10, p. 38-39]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be given by f=ghf=g-h, where $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable. Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial_{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
The next result is a combination of Theorem 2.2 with [10, Corollary on p. 52].
Theorem 2.3.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function and ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a closed and convex set. If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and all $x\in{{\cal C}}$. As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $u\in\partial_{c}h({x^{*}})$ and all $x\in{{\cal C}}$.
Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$. A continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ has gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$ if there exists a Lipschitz constant $L>0$ such that $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$. Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [6, Proposition A.24], see also[12, Lemma 2.4.2].
Proposition 2.4.
Let $g:\mathbb{R}^{n}\to\mathbb{R}$ be a differentiable with gradient $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, $x\in{{\cal C}}$, $v\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$. If $x+\lambda v\in{{\cal C}}$, then $g(x+\lambda v)\leq g(x)+\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}v\lambda+\frac{L}{2}\|v\|^{2}\lambda^{2}.$
Proposition 2.5.
The function $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex if and only if $h(y)\geq h(x)+\langle u,y-x\rangle$, for all $x,y\in\mathbb{R}^{n}$ and all $u\in\partial h(x)$.
Proposition 2.6 ([23, Proposition 6.2.1]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. Let $(x^{k})_{k\in{\mathbb{N}}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$. If $\lim_{k\to+\infty}x^{k}={\bar{x}}$ and $\lim_{k\to+\infty}u^{k}={\bar{u}}$, then ${\bar{u}}\in\partial h({\bar{x}})$.
Proposition 2.7 ([23, Proposition 6.2.2]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. The mapping $\partial h$ is locally bounded, i.e. the image $\partial h(B)$ of a bounded set $B\subset\mathbb{R}^{n}$ is a bounded set in $\mathbb{R}^{n}$.
In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [2, Lemma 13.13, Ch. 13, p. 387].
Lemma 2.8.
Let $(a_{k})_{k\in{\mathbb{N}}}$ and $(b_{k})_{k\in{\mathbb{N}}}$ be nonnegative sequences of real numbers satisfying
$a_{k+1}\leq a_{k}-b_{k}\beta_{k}+\frac{A}{2}\beta_{k}^{2},\qquad k=0,1,2,\ldots,$
where $\beta_{k}=2/(k+2)$ and $A$ is a positive number. Suppose that $a_{k}\leq b_{k}$, for all $k$. Then
$a_{k}\leq\frac{2A}{k}$, for all $k=1,2,\ldots.$
$\min_{\ell\in\{\lfloor\frac{k}{2}\rfloor+2,\ldots,k\}}b_{\ell}\leq\frac{8A}{k-2}$, for all $k=3,4,\ldots,$ where, $\lfloor k/2\rfloor=\max\left\{n\in\mathbb{N}:~n\leq k/2\right\}.$
We conclude this section by recalling the definitions of some norms that will be used later. For $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$ and $p\in[1,+\infty)$, the pp–norm is $\|x\|_{p}:=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{\!1/p}$. In particular, the $\ell_{1}$–norm is $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, and the $\ell_{\infty}$–norm is $\|x\|_{\infty}:=\max_{1\leq i\leq n}|x_{i}|$.
3 Piecewise-star-convex functions
In this section, we introduce the concept of piecewise-star-convexity, a relaxation of the classical star–convexity notion from [37]. Many DC functions that are nonconvex nevertheless satisfy this weaker property. As will be shown in a later section, by identifying and exploiting piecewise–star–convexity we can broaden the class of problems for which the classical Frank–Wolfe algorithm attains an ${\cal O}(1/k)$ convergence rate. We begin by recalling the definition of star–convexity from [37] and then present its relaxed counterpart, the piecewise-star-convex notion, which will be central to our analysis.
Definition 3.1.
Let ${{\cal C}}\subset{\mathbb{R}}^{n}$ be a convex set. A function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ is star–convex on $\cal C$}} if its set of global minimizers $X^{*}:=\arg\min_{{\cal C}}f$ is nonempty and, for any $x^{*}\in X^{*}$,
$f(\lambda x^{*}+(1-\lambda)x)\leq\lambda f(x^{*})+(1-\lambda)f(x),\qquad\forall\,x\in{{\cal C}},\ \forall\,\lambda\in[0,1].$ (1)
Every convex function with a nonempty set of global minimizers is star–convex, but the converse does not hold. We present two nonconvex star–convex functions from [37] (see also [35] for additional examples) to highlight that they are DC functions, each expressible as the difference of two convex functions.
Example 3.2.
The star–convex function $\phi(t)=|t|(1-e^{-|t|})$ is not convex. In addition, $\phi$ is a difference of two convex functions. Indeed, considering the following two convex functions $\varphi(t)=|t|(1-e^{-|t|})+e^{|t|}$ and $\psi(t)=e^{|t|}$, we have $\phi=\varphi-\psi$. Similarly, consider the star–convex function $f(s,t)=s^{2}t^{2}+s^{2}+t^{2}$. The function $f$ is not convex, but is a difference of two convex functions. Indeed, letting the following convex functions $g(s,t)=(s^{2}+t^{2})^{2}+s^{2}+t^{2}$ and $h(s,t)=s^{2}t^{2}+s^{4}+t^{4}$, we have f=ghf=g-h.
The requirement in (1) is formulated with respect to a fixed minimizer $x^{*}$ and must hold for all $x\in{\cal C}$. A milder, yet still powerful, relaxation is to allow the reference minimizer to depend on $x$. In many DC models, the reference minimizer that makes (1) valid for a given $x$ may vary with $x$, which is precisely what piecewise-star-convexity captures. Even this global condition can be too restrictive when $f$ exhibits multiple “basins” of favorable geometry. A natural remedy is to partition the domain into regions within which a local star–type inequality holds with respect to a (local) minimizer:
Definition 3.3.
Let ${\cal C}\subset\mathbb{R}^{n}$ be convex and let ${\cal P}=\{\,{\cal C}_{i}\,:\,i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets with pairwise disjoint interiors and $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$} with respect to $\cal P$} if, for every $i\in I$,
  • (i)
    the set of global minimizers of $f$ on ${\cal C}_{i}$, denoted $X_{i}^{*}$, is nonempty; and
  • (ii)
    for each $x\in{\cal C}_{i}$ there exists $x_{x}^{*}\in X_{i}^{*}$ such that
    $f\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x),\qquad\forall\,\lambda\in[0,1].$ (2)
Remark 1.
When the partition is trivial (${\cal P}=\{{\cal C}\}$) and $X_{1}^{*}=X^{*}$, Definition 3.3 reduces to Definition 3.1. In practice, piecewise-star-convexity often arises when $f$ is convex on each region ${\cal C}_{i}$ but not across different regions. In this setting, the minimizers $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ are global minimizers within their corresponding cells ${\cal C}_{i}$, although they may be only local (and possibly non–global) minimizers of $f$ on the whole domain $\cal C$}. This structure ensures that star–type inequalities hold on each cell (with respect to its minimizers $X_{i}^{*}$), even if such inequalities fail when crossing cell boundaries.
We emphasize that in Definition 3.3, the collection of subsets in item (ii) need not be finite: the piecewise-star-convexity property may still hold when $\cal C$} is covered by infinitely many convex regions. The following example illustrates that finiteness cannot, in general, be assumed.
Example 3.4.
Consider the set ${\cal C}=[0,1]$ and the infinite set $S:=\{0\}\cup\{1/k:k\in\mathbb{N}\}\subset[0,1],$ which is compact, infinite, and has an accumulation point at 0. Define $f:[0,1]\to\mathbb{R}$ by
$f(x)=\min_{c\in S}|x-c|^{2},\qquad x\in[0,1],$
see the Figure 1. The set of global minimizers is $X^{*}=\operatorname*{argmin}_{x\in[0,1]}f(x)=S$ with $\min f=0$. For each $x\in[0,1]$, choose $x_{x}^{*}\in S$ attaining the minimum in the definition of $f$, and let
$V_{x_{x}^{*}}:=\{\,y\in[0,1]:|y-x_{x}^{*}|\leq|y-c|\ \forall c\in S\,\}.$
These sets form a countable partition $0,1]=\bigcup_{k=0}^{\infty}V_{k}$, with
$V_{1}=\Bigl{[}\tfrac{3}{4},\,1\Bigr{]},\quad V_{k}=\Bigl{[}\tfrac{1/k+1/(k+1)}{2},\ \tfrac{1/k+1/(k-1)}{2}\Bigr{]},\ k\geq 2,\quad V_{0}=\{0\}.$
By construction, $x\in V_{x_{x}^{*}}$ and the whole segment $x_{x}^{*},x]$ lies in $V_{x_{x}^{*}}$. Moreover, $f(y)=|y-x_{x}^{*}|^{2}$ for all $y\in[x_{x}^{*},x]$. Hence, for any $\lambda\in[0,1]$,
$f\big{(}\lambda x_{x}^{*}+(1-\lambda)x\big{)}=\big{|}\lambda x_{x}^{*}+(1-\lambda)x-x_{x}^{*}\big{|}^{2}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x).$
Since $f(x_{x}^{*})=0$, the piecewise-star-convexity inequality holds. However, covering $\cal C$} by convex sets satisfying (i)–(ii) requires all sets $V_{k}$ above, so any such cover must be infinite.
[…]
Figure 1: Plot of the function $f(x)=\min_{c\in S}(x-c)^{2}$ for $S=\{0,\tfrac{1}{8},\tfrac{1}{7},\tfrac{1}{6},\tfrac{1}{5},\tfrac{1}{4},\tfrac{1}{3},\tfrac{1}{2},1\}$. The black segments indicate the active arc of the lower envelope, the dashed lines mark the boundaries of $V(x_{x}^{*})$, and the solid points on the $x$-axis represent the centers $c\in S$.
We next present a prototypical example of a function that is piecewise-star-convex. This model arises naturally in sparse estimation, robust quadratics, and regularized control. The $\ell_{1}$ penalty acts as a sign–dependent linear perturbation of a positive–definite quadratic. By partitioning $\mathbb{R}^{n}$ into orthants with fixed coordinate signs, the function reduces on each region to a strictly convex quadratic with a unique minimizer. These cell minimizers provide the reference points in Definition 3.3, thereby establishing piecewise-star-convexity even though, globally, such minimizers may be only local.
For use in the examples below and to simplify notation, define
$\{-1,1\}^{n}=\{s:=(s_{1},\ldots,s_{n})\in\mathbb{R}^{n}:~s_{i}\in\{-1,1\},~\forall i\in\{1,\ldots,n\}\}.$ (3)
Example 3.5.
Let $Q\in\mathbb{R}^{n\times n}$ be positive definite and let $\beta>0$. Consider $f:{\mathbb{R}}^{n}\to\mathbb{R}$ given by
$f(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$ (4)
For each $s\in\{-1,1\}^{n}$, define the closed orthant
$\Omega_{s}=\bigl{\{}\,x\in\mathbb{R}^{n}:~s_{i}x_{i}\geq 0,\,\text{for }i=1,\ldots,n\bigr{\}},\qquad\qquad\mathbb{R}^{n}=\bigcup_{s\in\{-1,1\}^{n}}\Omega_{s}.$
On each $\Omega_{s}$, the signs of the coordinates are fixed and $\|x\|_{1}=\sum_{i}|x_{i}|=\sum_{i}s_{i}x_{i}=s^{\top}x$, for $x\in\Omega_{s}$. Hence, the restriction of $f$ to $\Omega_{s}$ is the strictly convex quadratic
$f_{s}(x)=\tfrac{1}{2}\,x^{\top}Qx-\beta\,s^{\top}x,\qquad x\in\Omega_{s},$
with constant Hessian $Q$. Since $Q$ is positive definite, $f_{s}$ is strongly convex on $\mathbb{R}^{n}$ with modulus $\lambda_{\min}$, the minimizer $x_{s}^{*}$ is unique. Considering that $f$ coincides with the strictly convex quadratic $f_{s}$ on each orthant $\Omega_{s}$ and $x_{s}^{*}$ is the (unique) minimizer of $f$ on $\Omega_{s}$, the convexity inequality
$f\bigl{(}\lambda x_{s}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{s}^{*})+(1-\lambda)f(x),\qquad\forall\,x\in\Omega_{s},\ \forall\,\lambda\in[0,1],$
holds. Therefore $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the full partition $\{\Omega_{s}\}_{s\in\{-1,1\}^{n}}$.
Next we preste a constrained DC optimization problem motivated by contrastive learning formulations.
Example 3.6.
Given two disjoint sets of reference points, a set of positive examples $\mathcal{P}=\{\bar{p}_{1},\dots,\bar{p}_{m}\}\subset\mathbb{R}^{n}$ and a set of negative examples $\mathcal{N}=\{\bar{q}_{1},\dots,\bar{q}_{r}\}\subset\mathbb{R}^{n}$, the objective is to identify a point $x$ that lies close to the positive references while remaining far from the negative ones. Here all distances are measured with respect to the standard Euclidean metric. To remain consistent with our general framework, we impose that the search is carried out over a closed and convex feasible set $\mathcal{C}\subset\mathbb{R}^{n}$. Formally, we study the problem
$\min_{x\in\mathcal{C}}f(x),$ (5)
where the Euclidean contrastive objective is defined by
$f(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}-\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},\qquad x\in\mathbb{R}^{n},$ (6)
with $\lambda_{i}^{+}>0$ and $\lambda_{j}^{-}>0$ fixed weights. The squared $\ell_{2}$ terms pull $x$ toward the positive examples, while the negative $\ell_{1}$ term acts as a contrastive regularizer, promoting separation from the negative examples in a coordinate-dependent way. The objective (6) admits a natural DC decomposition of the form
$g(x):=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2},\qquad h(x):=\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1},$ (7)
where $g$ is smooth and strongly convex, while $h$ is convex but nondifferentiable. Hence similarly to Example 3.5 we can shoe that f=ghf=g-h belongs to the class of piecewise-star-convex DC models. In fact, let $m\geq 1$ and $\lambda_{i}^{+}>0$ for $i=1,\ldots,m$, and define $\Lambda^{+}:=\sum_{i=1}^{m}\lambda_{i}^{+}>0$. Consider $f:\mathbb{R}^{n}\to\mathbb{R}$ given in (6). Then $f$ is piecewise–star–convex on $\mathbb{R}^{n}$ with respect to the partition ${\cal P}=\{{\cal C}_{S}\}_{S}$ induced by the hyperplane arrangement $\{x_{k}=\bar{q}_{j,k}:~j=1,\ldots,r,\ k=1,\ldots,n\}$, where for $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$,
${\cal C}_{S}:=\bigl{\{}\,x\in\mathbb{R}^{n}:\ s_{j,k}\,(x_{k}-\bar{q}_{j,k})\geq 0,\ \forall j=1,\ldots,r,\ \forall k=1,\ldots,n\,\bigr{\}}.$
Fix $S=(s_{1},\ldots,s_{r})$ with $s_{j}\in\{-1,1\}^{n}$ and consider $x\in{\cal C}_{S}$. For every $j$ and $k$ we have $|x_{k}-\bar{q}_{j,k}|=s_{j,k}(x_{k}-\bar{q}_{j,k})$ on ${\cal C}_{S}$, hence
$\sum_{j=1}^{r}\lambda_{j}^{-}\,\|x-\bar{q}_{j}\|_{1}\;=\;\sum_{j=1}^{r}\lambda_{j}^{-}\sum_{k=1}^{n}s_{j,k}(x_{k}-\bar{q}_{j,k})\;=\;v_{S}^{\top}x-c_{S},$
where $v_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}\in\mathbb{R}^{n}$ and $c_{S}:=\sum_{j=1}^{r}\lambda_{j}^{-}\,s_{j}^{\top}\bar{q}_{j}$ is a constant on ${\cal C}_{S}$. Moreover,
$\sum_{i=1}^{m}\lambda_{i}^{+}\,\|x-\bar{p}_{i}\|_{2}^{2}\;=\;\Lambda^{+}\|x\|_{2}^{2}-2\,b^{\top}x\;+\;d,$
with $b:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\bar{p}_{i}$ and $d:=\sum_{i=1}^{m}\lambda_{i}^{+}\,\|\bar{p}_{i}\|_{2}^{2}$. Therefore, on ${\cal C}_{S}$ the function $f$ reduces to the strongly convex quadratic
$f_{S}(x)\;=\;\Lambda^{+}\|x\|_{2}^{2}-(2b+v_{S})^{\top}x\;+\;(d-c_{S}),$
whose Hessian is $2\Lambda^{+}I$ and is positive definite. Since ${\cal C}_{S}$ is closed and convex and $f_{S}$ is strongly convex, $f_{S}$ admits a unique minimizer $x_{S}^{\ast}\in{\cal C}_{S}$. By convexity of $f_{S}$ on the convex set ${\cal C}_{S}$, for all $x\in{\cal C}_{S}$ and all $\lambda\in[0,1]$,
$f\bigl{(}\lambda x_{S}^{\ast}+(1-\lambda)x\bigr{)}\;\leq\;\lambda f(x_{S}^{\ast})+(1-\lambda)f(x).$
This is precisely the star–type inequality required in Definition 3.3 with anchor $x_{S}^{\ast}$ on the cell ${\cal C}_{S}$. Since the interiors of the cells $\{{\cal C}_{S}\}$ are pairwise disjoint and $\bigcup_{S}{\cal C}_{S}=\mathbb{R}^{n}$, the claim follows.
The structure f=ghf=g-h reflects the central principle of contrastive learning: the quadratic component promotes similarity with positive references, while the $\ell_{1}$ component enforces dissimilarity from negative ones. For comprehensive discussions of contrastive learning in the Euclidean setting, see [24]. In the next proposition, we establish that, under a suitable dominance condition on the positive weights, the function $f$ in (6) is coercive and therefore attains at least one global minimizer of (5).
The next example shows how the minimum of convex quadratics induces a polyhedral partition of the space, a construction closely related to Voronoi diagrams and with applications in clustering, facility location, geometry of distance and constrained optimization (see [Libertietall2014] and [1] for a comprehensive survey on this subject). This construction is also connected to Example 3.4.
Example 3.7.
Let $Q\in\mathbb{R}^{n\times n}$ be positive definite, and let $b_{1},\dots,b_{m}\in\mathbb{R}^{n}$ and $c_{1},\dots,c_{m}\in\mathbb{R}$. Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
$f(x):=\min_{1\leq i\leq m}\varphi_{i}(x),\qquad\qquad\varphi_{i}(x):=\tfrac{1}{2}\,x^{\top}Qx+b_{i}^{\top}x+c_{i},\qquad i=1,\dots,m.$ (8)
Consider the partition ${\cal P}:=\{{\cal C}_{i}:~i=1,\ldots,m\}$ given by
${\cal C}_{i}=\bigl{\{}x\in\mathbb{R}^{n}:\ \varphi_{i}(x)\leq\varphi_{j}(x)\ \ \forall j=1,\dots,m\bigr{\}}=\bigcap_{j=1}^{m}\bigl{\{}x\in\mathbb{R}^{n}:\ (b_{i}-b_{j})^{\top}x\leq c_{j}-c_{i}\bigr{\}},$
since $\varphi_{i}(x)-\varphi_{j}(x)=(b_{i}-b_{j})^{\top}x+(c_{i}-c_{j})$. Hence each ${\cal C}_{i}$ is a closed convex polyhedron. By construction, we have
$f(x)=\varphi_{i}(x)\qquad\forall\,x\in{\cal C}_{i}.$
Because $Q$ is positive definite, each $\varphi_{i}$ is strongly convex; therefore $f$ coincides with a strongly convex quadratic on each cell ${\cal C}_{i}$. In particular, $f$ is convex on each ${\cal C}_{i}$, hence star–convex there, and thus piecewise star–convex with respect to the polyhedral partition $\{{\cal C}_{i}\}_{i=1}^{m}$.
Next, observe that $f$ admits a DC representation f=ghf=g-h with a smooth convex component $g$ and a convex (typically nonsmooth) component $h$. Indeed,
$\min_{1\leq i\leq m}\Bigl{\{}\tfrac{1}{2}\,x^{\top}Qx+b_{i}^{\top}x+c_{i}\Bigr{\}}=\tfrac{1}{2}\,x^{\top}Qx+\min_{1\leq i\leq m}\bigl{\{}b_{i}^{\top}x+c_{i}\bigr{\}}=\tfrac{1}{2}\,x^{\top}Qx-\max_{1\leq i\leq m}\bigl{\{}-b_{i}^{\top}x-c_{i}\bigr{\}}.$
Therefore, we have f(x)=g(x)h(x)f(x)=g(x)-h(x) with the components given by $g(x):=\tfrac{1}{2}\,x^{\top}Qx)$ and $h(x):=\max_{1\leq i\leq m}\bigl{(}-b_{i}^{\top}x-c_{i}\bigr{)}$. Clearly $g$ is $C^{\infty}$ and convex, while $h$ is the pointwise maximum of affine functions, hence convex, polyhedral, and Lipschitz.
A particularly relevant special case of (8) is obtained by taking pairwise distinct points $a_{1},\dots,a_{m}\in\mathbb{R}^{n}$ and setting $f(x)=\min_{1\leq i\leq m}\,\|x-a_{i}\|^{2},\qquad x\in\mathbb{R}^{n}.$
Next we show that a basic robustness property of piecewise–star–convexity is that the star–type inequality is preserved by adding a convex term.
Proposition 3.8.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, and let $g:\mathbb{R}^{n}\to\mathbb{R}$ be convex on $\cal C$}. Then $f+g$ is piecewise–star–convex on $\cal C$} with respect to $\cal P$}.
Proof.
Fix $i\in I$ and $x\in{\cal C}_{i}$. By Definition 3.3, there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ such that
$f\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda f(x_{x}^{*})+(1-\lambda)f(x),\quad\forall\,\lambda\in[0,1].$
Since $g$ is convex on ${\cal C}_{i}$,
$g\bigl{(}\lambda x_{x}^{*}+(1-\lambda)x\bigr{)}\leq\lambda g(x_{x}^{*})+(1-\lambda)g(x),\quad\forall\,\lambda\in[0,1].$
Summing the inequalities yields the star–type inequality for $f+g$ on ${\cal C}_{i}$ with the same base point $x_{x}^{*}$. As the argument holds for every $i\in I$ and $x\in{\cal C}_{i}$, the claim follows. ∎
Remark 2.
If your definition of piecewise–star–convexity explicitly requires the reference points to be minimizers of the function under consideration on each cell, then the statement above remains valid provided $(f+g)$ attains a (local) minimum on each ${\cal C}_{i}$ (e.g., when each ${\cal C}_{i}$ is compact or $(f+g)$ is coercive there). In that case one may choose the reeference points among $\arg\min_{{\cal C}_{i}}(f+g)$.
The next example combines Example 3.5 with Proposition 3.8 to show that a widely used optimization objective is piecewise–star–convex.
Example 3.9.
Let $A\in\mathbb{R}^{n\times n}$ be a full rank column, equivalently, $Q:=A^{\top}A$ a positive definite matrix, let $b\in\mathbb{R}^{n}$, and let $\beta>0$. Consider the function $f:{\mathbb{R}}^{n}\to\mathbb{R}$ defined by
$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{\,2}-\beta\|x\|_{1},\qquad x\in\mathbb{R}^{n}.$
Expanding the quadratic term gives the equivalent representation
$f(x)\;=\;\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}-c^{\top}x\;+\;\tfrac{1}{2}\|b\|_{2}^{2},\qquad Q:=A^{\top}A,\qquad c:=A^{\top}b.$ (9)
Equivalently, write $f=h+g$ with $h(x):=\tfrac{1}{2}\,x^{\top}Qx-\beta\|x\|_{1}$ and $g(x):=-\,c^{\top}x+\tfrac{1}{2}\|b\|_{2}^{2}$. By Example 3.5, $h$ is piecewise–star–convex on the orthant partition when $Q\succ 0$; since $g$ is convex (affine), Proposition 3.8 implies that $f=h+g$ is also piecewise–star–convex on the same partition.
Proposition 3.10.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{\mathcal{C}_{i}\}_{i\in I}$, where each $\mathcal{C}_{i}$ is closed and convex, interiors are pairwise disjoint, and $\bigcup_{i\in I}\mathcal{C}_{i}=\mathcal{C}$. Let $A\in\mathbb{R}^{n\times n}$ be invertible and $b\in\mathbb{R}^{n}$. Define the affine map
$\Psi(z):=Az+b,\qquad z\in\mathbb{R}^{n},$
the transformed domain and cells $\tilde{\mathcal{C}}:=A^{-1}(\mathcal{C}-b)$ and $\tilde{\mathcal{C}}_{i}:=A^{-1}(\mathcal{C}_{i}-b)$ and the function $\tilde{f}:\tilde{\mathcal{C}}\to\mathbb{R}$ by
$\tilde{f}(z):=f(Az+b),\qquad z\in\tilde{\mathcal{C}}$.
Then $\tilde{f}$ is piecewise–star–convex on $\tilde{\mathcal{C}}$ with respect to the partition $\tilde{\mathcal{P}}=\{\tilde{\mathcal{C}}_{i}\}_{i\in I}$.
Proof.
(i) Since $A$ is invertible, $\Psi$ is a bijective affine homeomorphism with inverse $\Psi^{-1}(x)=A^{-1}(x-b)$. If $z\in\tilde{\mathcal{C}}$ then $x:=\Psi(z)=Az+b\in\mathcal{C}$, so $\tilde{f}(z)=f(x)$ is well defined. Because $A^{-1}$ and the translation $x\mapsto x-b$ are continuous linear/affine maps, each $\tilde{\mathcal{C}}_{i}=A^{-1}(\mathcal{C}_{i}-b)$ is closed and convex; interiors remain pairwise disjoint and $\bigcup_{i}\tilde{\mathcal{C}}_{i}=\tilde{\mathcal{C}}$. Thus $\tilde{\mathcal{P}}$ is a partition of $\tilde{\mathcal{C}}$.
(ii) For each $i\in I$, let $X_{i}^{*}:=\arg\min_{\mathcal{C}_{i}}f$ and $\tilde{X}_{i}^{*}:=\arg\min_{\tilde{\mathcal{C}}_{i}}\tilde{f}$. We claim
$\tilde{X}_{i}^{*}=A^{-1}(X_{i}^{*}-b).$ (10)
Indeed, for $z\in\tilde{\mathcal{C}}_{i}$ write $x=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Then $\tilde{f}(z)=f(x)$ and hence
$\min_{z\in\tilde{\mathcal{C}}_{i}}\tilde{f}(z)=\min_{x\in\mathcal{C}_{i}}f(x),\qquad z\in\tilde{X}_{i}^{*}\ \Longleftrightarrow\ x=\Psi(z)\in X_{i}^{*}.$
Equivalently, $z=A^{-1}(x-b)$ with $x\in X_{i}^{*}$, proving (10).
(iii) Fix $i\in I$ and $z\in\tilde{\mathcal{C}}_{i}$, and set $x:=\Psi(z)=Az+b\in\mathcal{C}_{i}$. Because $f$ is piecewise–star–convex on $\mathcal{C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that, for all $\lambda\in[0,1]$,
$f\big{(}(1-\lambda)x+\lambda x_{x}^{*}\big{)}\leq(1-\lambda)f(x)+\lambda f(x_{x}^{*}).$ (11)
Define $z_{z}^{*}:=A^{-1}(x_{x}^{*}-b)$. By (10), $z_{z}^{*}\in\tilde{X}_{i}^{*}$. Using the affinity of $\Psi$,
$\Psi\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}=(1-\lambda)\Psi(z)+\lambda\Psi(z_{z}^{*})=(1-\lambda)x+\lambda x_{x}^{*}.$
Therefore,
$\tilde{f}\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}$ $=f\!\left(\Psi\big{(}(1-\lambda)z+\lambda z_{z}^{*}\big{)}\right)=f\big{(}(1-\lambda)x+\lambda x_{x}^{*}\big{)}$
$\leq(1-\lambda)f(x)+\lambda f(x_{x}^{*})=(1-\lambda)\tilde{f}(z)+\lambda\tilde{f}(z_{z}^{*}),$
where the inequality uses (11) and the identities $f(x)=\tilde{f}(z)$, $f(x_{x}^{*})=\tilde{f}(z_{z}^{*})$. Hence, for every $z\in\tilde{\mathcal{C}}_{i}$ there exists $z_{z}^{*}\in\tilde{X}_{i}^{*}$ such that the star–type inequality holds on $\tilde{\mathcal{C}}_{i}$. Since $i$ was arbitrary, $\tilde{f}$ is piecewise–star–convex on $\tilde{\mathcal{C}}$ with respect to $\tilde{\mathcal{P}}$. ∎
The next result establishes a first-order inequality that holds for any piecewise-star-convex. This property will play a key role in the convergence analysis of our algorithm.
Proposition 3.11.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz on a convex set $\cal C$}. Assume $f$ is piecewise–star–convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$ of Definition 3.3. Fix $i\in I$ and $x\in{\cal C}_{i}$. Then there exists $x_{x}^{*}\in X_{i}^{*}:=\arg\min_{{\cal C}_{i}}f$ (a minimizer on the cell ${\cal C}_{i}$; hence, a local minimizer for $f$ on $\cal C$}) such that
$f(x_{x}^{*})-f(x)\ \geq\ f^{\circ}(x;\,x_{x}^{*}-x)\ \geq\ v^{\top}(x_{x}^{*}-x)\qquad\forall\,v\in\partial_{c}f(x),$ (12)
where $f^{\circ}$ denotes Clarke’s generalized directional derivative and $\partial_{c}f(x)$ the Clarke subdifferential at $x$.
Proof.
By piecewise–star–convexity on ${\cal C}_{i}$, there exists $x_{x}^{*}\in X_{i}^{*}$ such that
$f\bigl{(}x+\lambda(x_{x}^{*}-x)\bigr{)}\ \leq\ (1-\lambda)f(x)+\lambda f(x_{x}^{*}),\qquad\forall\,\lambda\in[0,1].$
Subtract $f(x)$, divide by $\lambda>0$, and let $\lambda$ goes to 0 to obtain $f^{\circ}(x;\,x_{x}^{*}-x)\ \leq\ f(x_{x}^{*})-f(x).$ By the Clarke calculus (e.g., Theorem 2.1), $f^{\circ}(x;d)=\max_{v\in\partial_{c}f(x)}v^{\top}d$, hence $v^{\top}(x_{x}^{*}-x)\leq f^{\circ}(x;\,x_{x}^{*}-x)$ for all $v\in\partial_{c}f(x)$, which proves (12). ∎
The next result shows that, under piecewise-star-convexity, every Clarke stationary point of the constrained problem is in fact a minimizer of $f$ on the cell to which it belongs.
Corollary 3.12.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be locally Lipschitz and piecewise-star-convex on $\cal C$} with respect to the partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$. Suppose that $\bar{x}\in{\cal C}$ is a Clarke stationary point of the constrained problem $\min_{x\in{\cal C}}f(x),$ that is,
$v^{\top}(x-\bar{x})\ \geq\ 0\qquad\forall\,v\in\partial_{c}f(\bar{x}),\ \forall\,x\in{\cal C}.$ (13)
If, in addition, $\bar{x}\in{\cal C}_{i_{0}}$ for some $i_{0}\in I$, then $\bar{x}$ is a global minimizer of $f$ on the corresponding cell ${\cal C}_{i_{0}}$, namely, $\bar{x}\in X_{i_{0}}^{*}:=\arg\min_{x\in{\cal C}_{i_{0}}}f(x).$ In particular, this means that $f(\bar{x})=\min_{x\in{\cal C}_{i_{0}}}f(x).$
Proof.
Apply Proposition 3.11 with $x=\bar{x}$ and $i=i_{0}$ to obtain some $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ with
$f(x_{\bar{x}}^{*})-f(\bar{x})\ \geq\ v^{\top}(x_{\bar{x}}^{*}-\bar{x})\qquad\forall\,v\in\partial_{c}f(\bar{x}).$
By stationarity (13) (taking $x=x_{\bar{x}}^{*}\in{\cal C}_{i_{0}}\subset{\cal C}$), the right-hand side is $\geq 0$, hence $f(x_{\bar{x}}^{*})\geq f(\bar{x})$. Since $x_{\bar{x}}^{*}\in X_{i_{0}}^{*}$ minimizes $f$ on ${\cal C}_{i_{0}}$, we also have $f(x_{\bar{x}}^{*})\leq f(\bar{x})$. Therefore $f(x_{\bar{x}}^{*})=f(\bar{x})$ and $\bar{x}\in X_{i_{0}}^{*}$. ∎
Remark 3.
In the piecewise setting, $X_{i}^{*}$ consists of the global minimizers of $f$ restricted to the cell ${\cal C}_{i}$, which are in general only local minimizers on the whole set $\cal C$}. Corollary 3.12 refines the global statement for piecewise-star-convex functions by showing that Clarke stationarity implies cellwise optimality. Moreover, if the cell ${\cal C}_{i_{0}}$ containing $\bar{x}$ attains the global minimum value, i.e., $\min_{{\cal C}_{i_{0}}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}. This highlights the flexibility of piecewise-star-convexity: by allowing the domain to be partitioned into regions where the function behaves “locally convex,” one captures a broader class of DC objectives than with classical star–convexity, while still retaining descent inequalities sufficient to establish convergence of Frank–Wolfe variants.
4 The DC optimization problem
We are interested in solving the following constrained DC optimization problem
$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (14)
where ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact and convex set, $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function and $h:\mathbb{R}^{n}\to\mathbb{R}$ is convex function possibly non-differentiable, with domain on $\mathbb{R}^{n}$. Throughout the paper we assume that the gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$, i.e., there exists a Lipschitz constant $L>0$ such that
  • (A)
    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
Since we are assuming that ${\cal C}\subset{\mathbb{R}}^{n}$ is a compact set, its diameter is a finite number defined by
$\operatorname{diam}({\cal C}):=\max\left\{\|x-y\|:~x,y\in{\cal C}\right\}.$ (15)
Since ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact, the study of problem (14) is bounded from below. Then, the optimum value of the problem (14) satisfies $\infty<f^{*}:=\inf_{x\in{{\cal C}}}f(x)$ and optimal set $\cal C$$^{*}$ is non-empty. According to Theorem 2.3, the first-order optimality condition for problem (14) is stated as
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(x-{\bar{x}})\geq 0,\qquad\forall{\bar{u}}\in\partial g({\bar{x}}),\quad\forall x\in{{\cal C}}.$ (16)
In general, the condition (16) is necessary but not sufficient for optimality. A point ${\bar{x}}\in{{\cal C}}$ satisfying condition (16) is called a stationary point to problem (14). Consequently, all $x^{*}\in{{\cal C}}^{*}$ satisfies (16). We finish this section by presenting a variant of a classical example of DC optimization problem that fits the aforementioned requirements.
Example 4.1.
Let ${\cal C}_{i}\subset{\mathbb{R}}^{n}$ be an arbitrary set, for all $i=1,\ldots,m$. The square distance with respect to the set ${\cal C}_{i}$, denoted by $d^{2}_{{\cal C}_{i}}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is defined
$d^{2}_{{\cal C}_{i}}(x):=\inf_{y\in{\cal C}_{i}}\|x-y\|^{2}.$ (17)
In general, the function $d^{2}_{{\cal C}_{i}}$ is not convex. However, due to $\|x-y\|^{2}=\|x\|^{2}-2x^{T}y+\|y\|^{2}$, the distance function $d^{2}_{{\cal C}_{i}}$ becomes to
$d^{2}_{{\cal C}_{i}}(x):=\|x\|^{2}-\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right),$ (18)
which is a difference of two convex functions, see [23, Example 2.1.4]. Indeed, $d^{2}_{{\cal C}_{i}}=g-h$, where $g(x)=\|x\|^{2}$ is a convex quadratic function and $h(x)=\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)$ is a convex function given by the supremum of affine functions $x\mapsto\ell_{y}(x):=2x^{T}y-\|y\|^{2}$, for $y\in{\cal C}_{i}$. Let us state the constrained generalized Fermat-Weber location problem with square distances. For that, take $\omega_{i}\geq 0$, $i=1,\ldots,m$, such that $\sum_{i=1}^{m}\omega_{i}=1$ and note that
$\sum_{i=1}^{m}\omega_{i}d^{2}_{{\cal C}_{i}}(x)=\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right).$
Thus, for a given constrained set compact and convex ${\cal C}\subset{\mathbb{R}}^{n}$, the DC version of the constrained generalized Fermat-Weber location problem is stated as follows
$\begin{array}[]{c}\min_{x\in{{\cal C}}}\left(\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)\right).\end{array}$ (19)
Finally, observe that the objective function in problem (19) satisfies all the assumptions of problem 14. Moreover, in certain cases the objective function of the generalized Fermat–Weber problem (19) is weak–star–convex. For example, if $m=n$ and ${\cal C}_{i}=\{e_{i},-e_{i}\}$, where $e_{i}$ denotes the $i$-th canonical vector of $\mathbb{R}^{n}$, then the objective becomes $f(x)=\|x\|^{2}-2\sum_{i=1}^{n}\omega_{i}|x_{i}|+1,$ which, by the same reasoning as in Example 3.5, can be shown to be weak–star–convex.
5 Classical Frank–Wolfe algorithm with an adaptive stepsize
In this section, we present the classic Frank–Wolfe algorithm with an adaptive stepsize to solve problem (14) and provide its convergence analysis. Although we assume condition (A) for $g$, the Lipschitz constant will not be used explicitly to determine the stepsize. Instead, the stepsize is computed adaptively using the scheme introduced in [3, 5], which estimates the Lipschitz constant on the fly. Our convergence analysis shows that, for piecewise-star-convexfunctions, possibly non–differentiable and non–convex, the method achieves an ${\cal O}(1/k)$ convergence rate for both the function values and the duality gap. It is worth noting that a variant of the classic Frank–Wolfe algorithm for solving problem (14) was proposed in [26], where the stepsize depends on a curvature/Lipschitz–type constant of the gradient of $f$. For non–convex objective functions, the convergence analysis in [28] shows that such a variant attains only an ${\cal O}(1/\sqrt{k})$ rate in terms of reaching a stationary point.
Throughout this section, we assume access to a linear optimization oracle (LO oracle) that can minimize linear functions over the set $\cal C$}. The algorithm is stated below.
Algorithm 1.
Frank–WolfeC,f:=g-h algorithm
Step 0.
Select $x^{0}\in{{\cal C}}$ and $L_{0}>0$. Set $k=0$.
Step 1.
Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{0}\}$.
Step 2.
Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}(x_{k})$ as follows
$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}(x_{k}):=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (20)
Step 3.
If ${\omega}(x_{k})=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}(x_{k})|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (21)
Step 4.
If
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (22)
then set $j_{k}=j$ and go to Step 5. Otherwise, set $j=j+1$ and go to Step 3.
Step 5.
Set $\lambda_{k}:=\lambda_{j_{k}}$ and define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
$x^{k+1}:=x^{k}+\lambda_{k}(p^{k}-x^{k}),\qquad\quad L_{k+1}:=2^{j_{k}-1}L_{k}.$ (23)
Set $k\leftarrow k+1$, and go to Step 1.
Let us describe the main features of Algorithm 1. First of all, note that due to $x^{0}\in{C}$, $p^{k}\in{C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and C{C} be convex set, it follows from inductive arguments and first equality in (23) that $(x^{k})_{k\in{\mathbb{N}}}\subset{C}$. Note that (20) implies that
${\omega}(x_{k})\leq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall p\in{C}.$ (24)
Consequently, if ${\omega}(x_{k})=0$, then $(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0$, for all $p\in{C}$, i.e., $x^{k}$ is a stationary point to problem (14). Note that letting $p=x^{k}$ in (24) we conclude that
${\omega}(x_{k})\leq 0,\qquad k=0,1,\ldots.$ (25)
In order to simplify the notations, from now on we will use the following two notations:
${\omega}_{k}:={\omega}(x_{k}).$ (26)
Since $\omega_{k}=0$ implies that $x^{k}$ is a stationary point to problem (14). Thus, in view of (25) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$. The next result shows that the backtracking line–search in Algorithm 1 is well defined, identifies the smallest $j_{k}$ in Step 4, and establishes a descent inequality for the sequence $(x^{k})_{k\in{\mathbb{N}}}$.
Proposition 5.1.
For all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, the inequality (22) holds. Consequently, the number $j_{k}$ in Step 4 is well defined. Furthermore, $j_{k}$ is the smallest non-negative integer satisfying the following two conditions
$2^{j_{k}}L_{k}\geq 2L_{0},$ (27)
$f(x^{k}+\lambda_{k}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}_{k}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}.$ (28)
In addition, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. And the following inequality holds
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k},\qquad\qquad k=0,1,\ldots.$ (29)
Proof.
First we will prove that (22) holds for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$. It follows from Proposition 2.4 that
$g(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}.$ (30)
On the other hand, considering that $u^{k}\in\partial h(x^{k})$, it follows from Proposition (2.5) that
$h(x^{k}+\lambda_{j}(p^{k}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{k}-x^{k}\rangle.$ (31)
Thus, combining (30) with (31) and taking into account the equality in (20) and (25), some algebraic manipulations shows that
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega_{k}}|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L-{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2}.$
Hence, it follows from the last inequality that (22) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proofs the first statement. Consequently, $j_{k}$ is well defined and is the smallest non-negative integer satisfying (27) and (28). Therefore, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. We proceed with the proof of (29), by showing the following inequality
$|{\omega_{k}}|\lambda_{k}+\frac{{2^{j_{k}}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{k}^{2}\leq-\frac{1}{2}|{\omega_{k}}|\lambda_{k}.$ (32)
For that we separately consider two cases: $\lambda_{k}=|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$ and ${j_{k}}=1$. In the former, we have $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}=|{\omega_{k}}|/{\lambda_{k}}$,which substituting into the left hand side of (32) yields an equality. If now $\lambda_{k}=1$, then it follows from (21) that $\lambda_{k}=1\leq|{\omega_{k}}|/({2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2})$ or equivalently ${2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|$ and $\lambda_{k}=1$, which again substituting into the left hand side of (32) yields the inequality. Hence, (32) holds. Therefore, (29) follows by combining (28) with (32), which concludes the proof. ∎
For simplifying the notations we define the following constants
$\alpha:=2(L+L_{0})\operatorname{diam}({{\cal C}})^{2}>0.$ (33)
Lemma 5.2.
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{0}\leq L_{k}\leq L+L_{0},\quad\qquad k=0,1,\ldots.$ (34)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{\alpha}\right\},\quad\qquad k=0,1,\ldots.$ (35)
Proof.
We proceed by induction to prove (34). Since the Lipschitz constant $L>0$, (34) holds for $k=0$. Now, assume that (34) holds for some $k>0$. We separately consider two cases: $j_{k}=0$ and $j_{k}\geq 1$. First assume that $j_{k}=0$. In this case, Step 1 together with second equality in (23) imply that $L_{k+1}=2^{-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}=0$. Since we are under the assumption that (34) holds, we have $L_{k+1}=2^{-1}L_{k}<L_{k}\leq L+L_{0}$. Thus, we conclude that two inequalities in (34) holds for $k+1$ and $j_{k}=0$. Next, we assume that $j_{k}\geq 1$. Again, Step 1 together with second equality in (23) yield $L_{k+1}:=2^{j_{k}-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}\geq 1$. To prove the second one, we assume by contradiction that $L_{k+1}>L+L_{0}$. This condition and (23) imply that $2^{j_{k}-1}L_{k}>L$. Thus, the first part of Proposition 5.1 implies that $j=j_{k}-1$ satisfies (22). Since the condition $L_{k+1}>L+L_{0}$ together (23) also implies $2^{j_{k}-1}L_{k}>L_{0}$, we have a contradiction because Proposition 5.1 also implies that $j_{k}$ is the smallest non-negative integer satisfying (27) and (28). Therefore, (34) holds for $k+1$, which concludes the proof that (34) holds for all $k\in{\mathbb{N}}$. The inequality (35) is an immediate consequence of (15), (21), the second inequality in (34) and (33). ∎
Next, we state iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1. Since the analysis closely follows that in [4], the proofs are omitted.
Theorem 5.3.
$\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is a stationary point to problem (14). In addition, for every $N\in{\mathbb{N}}$, there holds
$\min\left\{|{\omega}_{k}|:~k=0,1,\ldots,N\right\}\leq\max\left\{{{2(f(x^{0})-f^{*})}},\sqrt{2\alpha{(f(x^{0})-f^{*})}}\right\}\frac{1}{\sqrt{N+1}}.$
Since $\cal C$}} is compact and the functions $g$ and $h$ are convex, we can define the following constants:
$\Theta:=\min\left\{\frac{2}{{\bar{\Gamma}}\operatorname{diam}({\cal C})},\frac{1}{\alpha}\right\}\qquad{\bar{\Gamma}}:=\max_{x\in{{\cal C}}}\{\|\nabla g(x)\|+\|u\|:~u\in\partial h(x)\},\qquad\beta:=\frac{{\bar{\Gamma}}^{2}}{{\Theta L_{0}}}.$
Since the proof of the next theorem follows a standard argument in the convergence analysis of line–search methods, we omit it.
Theorem 5.4.
Let $\epsilon>0$. Then, Algorithm 1 generates a point $x^{k}$ such that $|{\omega}_{k}|\leq\epsilon$, performing at most,
$\left(\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}\right)\left(2+\log_{2}\left({\beta}\frac{1}{\epsilon^{2}}\right)\right)={\cal O}(|\log_{2}(\epsilon)|\epsilon^{-2})$
evaluations of functions $f$, and $\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}+1={\cal O}(\epsilon^{-2})$ evaluations of gradients of $f$ and subgradients of $h$.
5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
In this section, we establish iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 when the objective function is a DC function f=ghf=g-h that is piecewise weak–star–convex, where the first component $g$ has a Lipschitz continuous gradient on ${\cal C}\subset{\mathbb{R}}^{n}$ and the second component $h$ is convex but not necessarily differentiable. As will become clear, the piecewise-star-convexity property plays a central role in our analysis. By combining this property with the descent estimates from Proposition 5.1, we obtain the following global complexity result for the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1.
Theorem 5.5.
Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a piecewise-star-convex on $\cal C$}, and let ${\cal P}=\{\,{\cal C}_{i}:i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets such that $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. For each $i\in I$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
$f_{i}^{*}:=\min_{x\in{\cal C}_{i}}f(x)$
be the corresponding minimum value. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exists $i_{0}\in I$ such that:
$f(x^{k})-f_{i_{0}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k}$, for all $k=1,2,\ldots.$
$\min_{\ell\in\left\{\lfloor\frac{k}{2}\rfloor+2,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-2}$, for all $k=3,4,\ldots,$ where $\lfloor k/2\rfloor=\max_{n\in\mathbb{N}}\left\{n\leq k/2\right\}.$
Proof.
By (28) in Proposition 5.1, we have
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega(x^{k})|\,\lambda_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (36)
On the other hand, by (21),
$\lambda_{k}\;=\;\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big{\{}-|\omega(x^{k})|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big{\}}.$
Hence, for any choice $\beta_{k}\in(0,1]$, the right-hand side of (36) at $\lambda=\lambda_{k}$ is bounded by its value at $\lambda=\beta_{k}$. Taking $\beta_{k}:=\tfrac{2}{k+2}$, we get
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
Using $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$, together with (23) and Lemma 5.2 (which yields $\tfrac{2^{\,j_{k}}L_{k}}{2}\leq L+L_{0}$), we obtain
$f(x^{k+1})\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;(L+L_{0})\,\operatorname{diam}({\cal C})^{2}\,\beta_{k}^{2}.$ (37)
Since $\cal C$} is compact, the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1 is bounded. Let $\bar{x}$ be a limit point, and let $(x^{k_{\ell}})_{\ell\in\mathbb{N}}$ be a subsequence such that $\lim_{\ell\to+\infty}x^{k_{\ell}}={\bar{x}}$. Because ${\cal P}=\{{\cal C}_{i}:i\in I\}$ is a partition of $\cal C$}, there exists $i_{0}\in I$ with $\bar{x}\in{\cal C}_{i_{0}}$. In addition, Theorem 5.3 implies that $\lim_{k\to+\infty}{\omega}_{k}=0$ and $\bar{x}$} is a stationary point to problem (14). Hence, by (3.12), $\bar{x}$ is a minimizer of $f$ on ${\cal C}_{i_{0}}$. Let $f_{i_{0}}^{*}$ denote the minimum value of $f$ on ${\cal C}_{i_{0}}$. Thus, $\lim_{\ell\to+\infty}f(x^{k_{\ell}})=f({\bar{x}})=f_{i_{0}}^{*}$. Moreover, (29) in Proposition 5.1 implies that $\big{(}f(x^{k})\big{)}_{k\in\mathbb{N}}$ is monotone nonincreasing. Therefore,
$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{0}}^{*},\qquad\text{and}\qquad f_{i_{0}}^{*}\leq f(x^{k})\quad\forall\,k\in\mathbb{N}.$ (38)
Since $f$ is piecewise weak–star–convex on $\cal C$}, Proposition 3.11 and Theorem 2.2 ensure that, for any $x^{\star}\in X_{i_{0}}^{*}$,
$f_{i_{0}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$
where $u^{k}\in\partial h(x^{k})$. Using (20) and (26), the definition of $\omega_{k}$ gives $0\geq f_{i_{0}}^{*}-f(x^{k})\geq\omega_{k},$ hence
$0\leq f(x^{k})-f_{i_{0}}^{*}\leq|\omega_{k}|\qquad\forall\,k\in\mathbb{N}.$ (39)
Now apply Lemma 2.8 to the sequences
$a_{k}=f(x^{k})-f_{i_{0}}^{*}\leq b_{k}=|\omega_{k}|,\qquad\beta_{k}=\frac{2}{k+2},\qquad A=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$
The recursion (37) together with (39) satisfies the hypotheses of Lemma 2.8, yielding
$f(x^{k})-f_{i_{0}}^{*}\leq\frac{4(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k},\qquad\forall k\geq 1,$
and
$\min_{\ell\in\{\lfloor k/2\rfloor+2,\ldots,k\}}|\omega_{\ell}|\leq\frac{16(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k-2}.\qquad\forall k\geq 3.$
These are precisely items (i)–(ii) of Theorem 5.5. ∎
Theorem 5.5 shows that, under piecewise–star–convexity, the exact–gradient Frank–Wolfe method with our stepsize achieves the convex-case rate: $f(x_{k})-f^{\star}=\mathcal{O}(1/k)$ and the optimality certificate $|\omega_{k}|$ decays at the same order. In contrast, the next section focuses on the finite–difference variant. We do not pursue complexity bounds there; instead we establish asymptotic convergence under mild consistency assumptions on the difference radius and the adaptive stepsizes: the gradient surrogate converges to $\nabla g$ along the iterates, the scheme is well posed, and every accumulation point is first-order stationary (equivalently, the Frank–Wolfe gap and $|\omega_{k}|$ vanish). Thus the Frank–Wolfe method provides projection-free, asymptotic guarantees compatible with the piecewise–star–convex framework, even without claiming the convex-case $\mathcal{O}(1/k)$ rate.
6 Frank–Wolfe algorithm with finite difference
In this section we introduce a projection-free Frank–Wolfe variant that replaces $\nabla g$ by a finite-difference surrogate with an adaptive stepsize, while retaining the linear minimization oracle of Algorithm 1. The scheme is designed to the DC model f=ghf=g-h from earlier sections, $g$ smooth convex and $h$ convex (via subgradients), and fits the DC piecewise–star–convexity framework developed in previous sections. Our analysis targets asymptotic convergence of the finite-difference method: under mild consistency conditions on the probing radii and stepsizes, the surrogate tracks $\nabla g$, the algorithm is well posed, and every accumulation point is first-order stationary (the Frank–Wolfe gap and $|\omega_{k}|$ vanish). We do not claim rates for this variant; Theorem 5.5 remains the $\mathcal{O}(1/k)$ benchmark for the exact-gradient case. Since piecewise–star–convexity objectives are convex within cells, stabilization in a cell yields local convex geometry, which motivates our stepsize choice.
Let $\{e_{i}\}_{i=1}^{n}$ denote the canonical basis of $\mathbb{R}^{n}$. For a differentiable $g:\mathbb{R}^{n}\to\mathbb{R}$ and a stepsize $s>0$, define the (componentwise) forward finite-difference approximation of $\nabla g(x)$ by
$D(x,s):=\bigl{(}D_{1}(x,s),\ldots,D_{n}(x,s)\bigr{)},\qquad D_{i}(x,s):=\frac{g(x+se_{i})-g(x)}{s},\quad i=1,\ldots,n.$ (40)
The next lemma (see [38, (8.4), p. 195]) provides an upper bound on the error of approximating $\nabla g(x)$ by $D(x,s)$.
Lemma 6.1.
Assume that $g$ satisfies the assumption (A). Then, there holds
$\|\nabla g(x)-D(x,s)\|\leq\frac{{\sqrt{n}L}}{2}s,\qquad\forall x\in\mathbb{R}^{n},\quad\forall s>0.$
Proof.
Setting $\nabla g(x)=(\partial_{1}g(x),\ldots,\partial_{n}g(x))$ and using (40) we have
$\|\nabla g(x)-D(x,s)\|^{2}\leq\sum_{i=1}^{n}\left(\partial_{i}g(x)-D_{i}(x,s)\right)^{2}.$ (41)
Applying Proposition 2.4 with $v=e_{i}$ and $\lambda=s$, we conclude that $D_{i}(x,s)-\partial_{i}g(x)\leq(Ls)/2$. Thus, (41) implies that $\|\nabla g(x)-D(x,s)\|^{2}\leq n((Ls)/2)^{2}$, which implies the desired inequality. ∎
The statement of the Frank–Wolfe algorithm to solve problem (14) with finite difference approximating the gradient $\nabla g$ is as follows:
Algorithm 2.
Frank–WolfeC,f:=g-h algorithm with finite difference
Step 0.
Select $x^{0},x^{1}\in{{\cal C}}$, $x^{0}\neq x^{1}$ and $L_{1}>0$. Set $k=1$.
Step 1.
Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $i=0$.
Step 2.
Set
$s_{ij}=\frac{2L_{1}}{2^{i+j}L_{k}\sqrt{n}}\|x^{k}-x^{k-1}\|.$ (42)
Step 3.
Use an “LO oracle” to compute an optimal solution $p^{ij}$ and the optimal value ${\omega}_{ij}(x_{k})$ as follows
$p^{ij}\in{\rm argmin}_{p\in{{\cal C}}}(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{ij}(x_{k}):=(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{ij}-x^{k}).$ (43)
Step 4.
If ${\omega}_{ij}(x_{k})<0$, then set $i_{k}=i$ and go to Step 5. Otherwise, set $i\leftarrow i+1$, and go to Step 2.
Step 5.
Compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{i_{k}j}(x_{k})|}{{2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{i_{k}j}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda^{2}\right\}.$ (44)
Step 6.
If
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}-\frac{2^{j}L_{k}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$ (45)
then set $j_{k}=j$ and go to Step 7. Otherwise, set $j\leftarrow j+1$ and go to Step 2.
Step 7.
Set $\lambda_{k}:=\lambda_{j_{k}}$, $p^{k}:=p^{i_{k}j_{k}}$, define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
$x^{k+1}:=x^{k}+\lambda_{k}(p^{k}-x^{k}),\qquad\quad L_{k+1}:=2^{j_{k}-1}L_{k}.$ (46)
Set $k\leftarrow k+1$, and go to Step 1.
Let us describe the main features of Algorithm 2. First of all, note that due to $x^{0},x^{1}\in{\cal C}$, $p^{k}\in{\cal C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and $\cal C$} be convex set, it follows from inductive arguments and first equality in (46) that $(x^{k})_{k\in{\mathbb{R}}}\subset{\cal C}$. Note that (43) implies that
${\omega}_{ij}(x_{k})\leq D(x^{k},s_{ij})^{T}(p-x^{k}),\qquad p\in{\cal C}.$ (47)
Hence, considering that $x^{k}\in{\cal C}$, letting $p=x^{k}$ in the last inequality we conclude that
${\omega}_{ij}(x_{k})\leq 0.$ (48)
In order to simplify the notations, from now on we will use the following two notations:
$s_{k}:=s_{i_{k}j_{k}},\qquad\qquad{\omega}_{k}:={\omega}_{i_{k}j_{k}}(x_{k}).$ (49)
Proposition 6.2.
Assume that in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$. Then,
$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0,\qquad\forall p\in{{\cal C}},$ (50)
i.e., $x^{k}$ is a stationary point to problem (14).
Proof.
If in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$, then it follows from (47) that
$0\leq(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall i\in{\mathbb{N}},\quad\forall p\in{{\cal C}}.$ (51)
Since (42) implies that $\lim_{i\to+\infty}s_{ij}=0$, it follows from (40) that $\lim_{i\to+\infty}D(x^{k},s_{ij})=\nabla g(x^{k})$. Hence, taking limit in (51) we conclude that (50) holds. ∎
In view of Proposition 6.2, (48) and (49) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$.
6.1 Well-definedness
Herein, we establish that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined.
Lemma 6.3.
Let $s_{ij}$ be define in (42) and assume that ${2^{j}L_{k}}\geq 2L$. Then, there holds
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{L_{1}}}{2}\|x^{k}-x^{k-1}\|.$
Proof.
By applying Lemma 6.1 with $x=x^{k}$ and $s=s_{ij}$ we obtain that
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{\sqrt{n}L}}{2}s_{ij}$
Therefore, using (42) and considering that ${2^{j}L_{k}}\geq 2L$ the desired inequality follows. ∎
Lemma 6.4.
Let $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$. Then, (45) holds and $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying the following two conditions $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined and satisfies the following inequality:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$ (52)
Proof.
It follows from Proposition 2.4 that
$g(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (53)
On the other hand, considering that $u^{k}\in\partial h(x^{k})$ it follows from Proposition (2.5) that
$h(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{i_{k}j}-x^{k}\rangle$ (54)
Thus, combining (53) with (54) and taking into account the equality in (43) and (48), some algebraic manipulations shows that
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (55)
To proceed with the proof we need first to prove that for $\lambda_{j}$ defined in (44) we have
$|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}\leq-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}.$ (56)
For that we separately consider two cases: $\lambda_{j}=|{\omega}_{i_{k}j}(x_{k})|/{(2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2})}$ and $\lambda_{j}=1$. In the former, we have $2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2}=|{\omega}_{i_{k}j}(x_{k})|/{\lambda_{j}}$,which substituting into the left hand side of (56) yields an equality. If now $\lambda_{j}=1$, then it follows from (44) that $\lambda_{j}=1\leq|{\omega}_{i_{k}j}(x_{k})|/({2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2})$ or equivalently ${2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}\leq|{\omega}_{i_{k}j}(x_{k})|$ and $\lambda_{j}=1$, which again substituting into the left hand side of (56) yields the inequality. Hence, (56) holds. Therefore, (55) becomes
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (57)
Since ${2^{j}L_{k}}\geq 2(L+L_{1}/2)>L$, we can apply Lemma 6.3 to obtain that
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}.$
Hence, considering that $2\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}\leq\|x^{k}-x^{k-1}\|^{2}+\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}$, we have
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Thus, combining the last inequality with (57) and considering that $\lambda_{j}<1$, we conclude that
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{L+L_{1}/2-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Considering that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$, the last inequality implies that (45) holds. Since the first trial $j$ in (45) is defined in Step 1, i.e., $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $2^{j+1}L_{k}\geq 2^{j}L_{k}$ we conclude that $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, $(x^{k})_{k\in{\mathbb{N}}}$ is well defined, which proof the first three statements of the lemma. Therefore, (52) follows from the definition of $j_{k}$ in Step 6, (45) and (46), which concludes the proof. ∎
Since the proof of the next lemma is similar to Lemma 5.2 it will be omitted.
Lemma 6.5.
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{1}\leq L_{k}\leq 2(L+L_{1}),\quad\qquad k=1,2,\ldots.$ (58)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}}\right\},\quad\qquad k=1,2,\ldots.$ (59)
6.2 Convergence analysis
In this section we analyze the convergence properties of the sequence $(x^{k})_{k\in{\mathbb{N}}}$. For that, we assume that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is infinite.
Lemma 6.6.
The sequence $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}}$ is monotone and non-increasing. Moreover, there holds:
$\lim_{k\to+\infty}\|x^{k}-x^{k-1}\|=0.$ (60)
As a consequence, $\lim_{k\to+\infty}f(x^{k})={\bar{f}}$ for some ${\bar{f}}\geq f^{*}$.
Proof.
Since $L_{k}\geq L_{1}$, for all $k\in{\mathbb{N}}$, it follows from Lemma 6.4 that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$
which, taking into account that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, yields
$f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\leq f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots..$
which implies the first statement of the lemma. We proceed to prove the second statement. For that we first note that Lemma 6.5 implies that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$. Thus, using again that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, it follows from (52) that
$\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\leq f(x^{k})-f(x^{k+1}),\qquad k=1,2,\ldots.$ (61)
Let $\ell\in{\mathbb{N}}$ be such that $\ell>1$. Summing up the inequality (61) from $1$ to $\ell$ we obtain
$\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}\leq f(x^{1})-f^{*}.$ (62)
Considering that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$, some algebraic manipulations show that
$\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}+\frac{L_{1}}{4}\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}.$
Combining (62) with the last inequality we obtain that
$\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\frac{4}{L_{1}}\Big{(}f(x^{1})-f^{*}+\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}\Big{)}.$
Since the last inequality holds for all $\ell\in{\mathbb{N}}$, we have $\sum_{k=1}^{+\infty}\|x^{k+1}-x^{k}\|^{2}\leq+\infty$. Thus, (60) holds. To prove the last statement, we first note that due to $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in\mathbb{N}}$ be a non-increasing sequence and bounded from below by $f^{*}$, it must be convergent. Thus, we set ${\bar{f}}=\lim_{k\to+\infty}\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)$. Hence, considering that
$f(x^{k})=f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}-({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\geq f^{*},$
by using (60) we conclude that $\lim_{k\to+\infty}f(x^{k})={\bar{f}}\geq f^{*}$ and the proof is concluded. ∎
For simplifying the notations we define the following constants
${\bar{\alpha}}:=2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}>0.$ (63)
Lemma 6.7.
The following inequality holds:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},$ (64)
for all $k=1,2,\ldots$.
Proof.
Combining inequalities (52) with (58) and (59), and also using (63), we conclude that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\min\left\{1,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$
for all $k=1,2,\ldots$. Thus, (64) is an immediate consequence of the previous inequality. ∎
In the next theorem we will use the following constant
$\Gamma:=2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}.$ (65)
Theorem 6.8.
The following statements hold:
  • (i)
    $\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is a stationary point to problem (14);
  • (ii)
    for every $N\in{\mathbb{N}}$, there holds
    $\min\left\{|{\omega}_{k}|:~k=1,\ldots,N\right\}\leq\max\left\{{\Gamma},\sqrt{{\bar{\alpha}}\Gamma}\right\}\frac{1}{\sqrt{N}}.$
Proof.
To prove the item (i), we first note that (64) is equivalent to
$\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2}\Big{)}-\Big{(}f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\Big{)},$ (66)
for all $k=1,2,\ldots$. It follows from Lemma 6.6 that $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}^{*}}$ monotone and non-increasing. Thus, considering that it is also bounded from below by $f^{*}$, we conclude that it converges. Hence, taking limit in (66) we have
$\lim_{k\to+\infty}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}=0,$
which implies the first statement of item (i). For proving the second statement of item (i), take $\bar{x}$} a limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ and subsequence $(x^{k_{i}})_{i\in{\mathbb{N}}}$ such that $\lim_{i\to+\infty}x^{k_{i}}={\bar{x}}$. Since $\cal C$}} is a compact set the sequence $(u^{k})_{k\in{\mathbb{N}}^{*}}$ is bounded. Thus, without loss of generality we assume that $\lim_{i\to+\infty}u^{k_{i}}={\bar{u}}$. On the other hand, Lemma 6.6 together with (42) imply that
$\lim_{k\to+\infty}s_{k}=0.$ (67)
Let $(s_{k_{i}})_{i\in{\mathbb{N}}^{*}}$ subsequence of $(s_{k})_{k\in{\mathbb{N}}^{*}}$. Thus, we conclude from (40) that
$\lim_{i\to+\infty}D(x^{k_{i}},s_{k_{i}})=\nabla g({\bar{x}}).$ (68)
Finally, using (43) we conclude that
$(D(x^{k_{i}},s_{k_{i}})-u^{k_{i}})^{T}(p-x^{k_{i}})\geq{\omega}_{k_{i}},\qquad\forall p\in{{\cal C}}.$
Therefore, taking limite in the last inequality and using (67) together (68) we obtain that
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(p-{\bar{x}})\geq 0,\qquad{\bar{u}}\in\partial g({\bar{x}}),\qquad\forall~x\in{{\cal C}},$
which implies that $\bar{x}$} is a stationary point to problem (14). To prove the item (ii), take $N\in{\mathbb{N}}^{*}$. Thus, (66) implies that
$\sum_{k=1}^{N}\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{1})+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\Big{)}-\Big{(}f(x^{N+1})+\frac{L_{1}}{2}\|x^{N}-x^{N-1}\|^{2}\Big{)}.$
Hence, we conclude that
$\min\left\{\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}:~k=1,\ldots N\right\}\leq\frac{1}{N}\left(f(x^{1})-f^{*}+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\right).$
It follows from the last inequality that there exists ${\bar{k}}\in\{1,\ldots N\}$ such that
$|{\omega}_{\bar{k}}|\leq\frac{1}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right),\qquad|{\omega}_{\bar{k}}|\leq\sqrt{\frac{{\bar{\alpha}}}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right)},$
which, by using (65), implies the item (ii). ∎
7 Conclusions
We analyzed Algorithm 1 and its finite–difference variant Algorithm 2 within the DC piecewise-star-convex framework. For piecewise-star-convex objectives, the exact–gradient method attains the convex–case rate $\mathcal{O}(1/k)$ for both the function suboptimality $f(x_{k})-f^{\star}$ and the optimality certificate (duality/Frank–Wolfe gap). For the finite–difference method, we established asymptotic convergence and an $\mathcal{O}(1/\sqrt{k})$ decay of the gap. A full iteration–complexity theory for the finite–difference variant under piecewise-star-convex remains open. Promising directions include: (i) matching $\mathcal{O}(1/k)$ rates via tighter control of finite–difference bias and adaptive radii, (ii) guarantees that account for possible transitions across piecewise-star-convex cells without stabilization assumptions, and (iii) extensions to stochastic oracles, away–steps, and line–search variants.
Funding
The first and third authors were supported by the Australian Research Council (ARC), Solving hard Chebyshev approximation problems through nonsmooth analysis (Discovery Project DP180100602). The second author was supported in part by CNPq - Brazil Grants 304666/2021-1 and Fapesc - TR nº 2024TR002238. This work was partially conducted while the second author visited the first and third authors at Deakin University in November 2022, with thanks to the host institution for funding and providing a stimulating scientific environment.